C. (Theory) • Prove that if X Exp(x) for some > 0, ² = Var(x) = 1 / 2
Answer #1

SOL XN Expld) ; do 999999 Then, flx; d)= de-dx ; x 3,0 V(x)= 82 = E[x²] - (E[x]) ² Now, Elx]= ddy -dx 2 & juxede de a-1 => E[up (kort zx J = (2x 17 8 1x de dix 3-1 - 3 - 2 e para W 2) 12 U = var Lx Z - to 102= var (x)2